# AC-PKAN: Attention-Enhanced and Chebyshev Polynomial-Based Physics-Informed Kolmogorov–Arnold Networks

### v1

@@ Abstract
Kolmogorov–Arnold Networks (KANs) have recently shown promise for solving partial differential equations (PDEs). Yet their original formulation is computationally and memory intensive, motivating the introduction of Chebyshev Type-I-based KANs (Chebyshev1KANs). Although Chebyshev1KANs have outperformed the vanilla KANs architecture, our rigorous theoretical analysis reveals that they still suffer from rank collapse, ultimately limiting their expressive capacity. To overcome these limitations, we enhance Chebyshev1KANs by integrating wavelet-activated MLPs with learnable parameters and an internal attention mechanism. We prove that this design preserves a full-rank Jacobian and is capable of approximating solutions to PDEs of arbitrary order. Furthermore, to alleviate the loss instability and imbalance introduced by the Chebyshev polynomial basis, we externally incorporate a Residual Gradient Attention (RGA) mechanism that dynamically re-weights individual loss terms according to their gradient norms and residual magnitudes. By jointly leveraging internal and external attention, we present AC-PKAN, a novel architecture that constitutes an enhancement to weakly supervised Physics-Informed Neural Networks (PINNs) and extends the expressive power of KANs. Experimental results from nine benchmark tasks across three domains show that AC-PKAN consistently outperforms or matches state-of-the-art models such as PINNsFormer, establishing it as a highly effective tool for solving complex real-world engineering problems in zero-data or data-sparse regimes. The code will be made publicly available upon acceptance.
@@ 1 Introduction
Numerical solutions of partial differential equations (PDEs) are essential in science and engineering (zienkiewicz2005finite, ; liu2009meshfree, ; fornberg1998practical, ; brebbia2012boundary, ). Physics-informed neural networks (PINNs) (lagaris1998artificial, ; raissi2019physics, ) have emerged as a promising approach in scientific machine learning (SciML), especially when data are unavailable or scarce. Traditional PINNs typically employ multilayer perceptrons (MLPs) (cybenko1989approximation, ) due to their ability to approximate nonlinear functions (hornik1989multilayer, ) and their success in various PDE-solving applications (yu2018deep, ; han2018solving, ).
However, PINNs encounter limitations, including difficulties with multi-scale phenomena (kharazmi2021hp, ), the curse of dimensionality in high-dimensional spaces (jagtap2020extended, ), and challenges with nonlinear PDEs (yuan2022pinn, ). These issues arise from both the complexity of PDEs and limitations in PINN architectures and training methods.
@@
To address these challenges, existing methods focus on improving both the internal architecture of PINNs and their external learning strategies. Internal improvements include novel architectures like Quadratic Residual Networks (Qres) (bu2021quadratic, ), First-Layer Sine (FLS) (wong2022learning, ), and PINNsformer (zhao2023pinnsformer, ). External strategies are discussed in detail in Section 2. Nevertheless, traditional PINNs based on MLPs still suffer from issues like lack of interpretability (cranmer2023interpretable, ), overfitting, vanishing or exploding gradients, and scalability problems (bachmann2024scaling, ). As an alternative, Kolmogorov–Arnold Networks (KANs) (liu2024kan, ), inspired by the Kolmogorov–Arnold representation theorem (kolmogorov1961representation, ; braun2009constructive, ), have been proposed to offer greater accuracy and interpretability. KANs can be viewed as a combination of Kolmogorov networks and MLPs with learnable activation functions (koppen2002training, ; sprecher2002space, ). Various KAN variants have emerged by replacing the B-spline functions (ss2024chebyshev, ; bozorgasl2024wav, ; xu2024fourierkan, ). Although they still face challenges (yu2024kan, ), KANs have shown promise in addressing issues like interpretability (liu2024kan2, ) and catastrophic forgetting (vaca2024kolmogorov, ) in learning tasks (samadi2024smooth, ). Recent architectures like KINN (wang2024kolmogorov, ) and DeepOKAN (abueidda2024deepokan, ) have applied KANs to PDE solving with promising results.
@@
Despite the potential of KANs, the original KAN suffers from high memory consumption and long training times due to the use of B-spline functions (shukla2024comprehensive, ). To address these limitations, we propose the Attention-Enhanced and Chebyshev Polynomial-Based Physics-Informed Kolmogorov–Arnold Networks (AC-PKAN). Our approach replaces B-spline functions with first-kind Chebyshev polynomials, forming the Cheby1KAN layer (chebykan2024, ), eliminating the need for grid storage and updates. Nevertheless, networks composed solely of stacked Cheby1KAN layers exhibit pronounced rank diminution (feng2022rank, ). By integrating Cheby1KAN with linear layers and incorporating internal attention mechanisms derived from input features, AC-PKAN addresses these limitations while efficiently modeling complex nonlinear functions and selectively emphasizing distinct aspects of the input features at each layer. Additionally, we introduce an external attention mechanism that adaptively reweights loss terms according to both gradient norms and point-wise residuals, thereby counteracting the large polynomial expansions and gradient magnitudes inherent in Cheby1KAN, mitigating residual imbalance and gradient flow stiffness, and ultimately enhancing training stability and efficiency. To our knowledge, AC-PKAN is the first PINN framework to integrate internal and external attention mechanisms into KAN layers, effectively addressing many issues of original KANs and PINNs. Our key contributions can be summarized as follows:
@@ 2 Related Works
External Learning Strategies for PINNs. Various external strategies have been proposed to address the limitations of PINNs. Loss weighting methods, such as PINN-LRA (wang2021understanding, ), PINN-NTK (wang2022and, ), and PINN-RBA (anagnostopoulos2024residual, ), rebalance loss terms using gradient norms, neural tangent kernels, and residual information to enhance training efficiency. Optimizer improvements like MultiAdam (yao2023multiadam, ) aid convergence during multi-scale training. Advanced sampling strategies, including AAS (tang2023adversarial, ), which combines optimal transport theory with adversarial methods, RoPINN (pan2024ro, ), which utilizes Monte Carlo sampling for regional optimization, RAR (wu2023comprehensive, ), which applies residual-driven resampling, and PINNACLE (lau2024pinnacle, ), which adaptively co-optimizes the selection of all types of training points, have been developed to improve performance. Enhanced loss functions like gPINN (yu2022gradient, ) and vPINN (kharazmi2019variational, ) incorporate gradient enhancement and variational forms, respectively. Adaptive activation functions in LAAF (jagtap2020locally, ) and GAAF (jagtap2020adaptive, ) accelerate convergence and handle complex geometries. Domain decomposition methods such as FBPINN (moseley2023finite, ) and hp-VPINN (kharazmi2021hp, ) train subnetworks on subdomains and use higher-order polynomial projections for refinement.
Variants of KAN. Since the seminal KAN (liu2024kan, ), numerous basis-function substitutions have been proposed to sharpen the speed–accuracy trade-off. FastKAN replaces cubic B-splines with radial basis functions (RBFs) for faster inference (li2024fastkan, ); Chebyshev1KAN and Chebyshev2KAN leverage first- and second-kind Chebyshev polynomials (chebykan2024, ; ss2024chebyshev, ); rKAN and fKAN introduce trainable rational- and fractional-orthogonal Jacobi bases (afzalaghaei2024rkan, ; afzalaghaei2024fkan, ); and FourierKAN substitutes spline coefficients with one-dimensional Fourier modes (fourierkan_github, ). A recent survey situates these derivatives in the broader landscape of Kolmogorov-inspired approximators (guilhoto2024deep, ). Preliminary benchmarks still crown Cheby1KAN as the current speed–accuracy frontrunner (kan_benchmarking_github, ).
@@ 3 Motivation and Methodology
$\begin{gathered}\mathcal{D}[u(\boldsymbol{x},t)]=f(\boldsymbol{x},t),\quad(\boldsymbol{x},t)\in\Omega,\\ \mathcal{B}[u(\boldsymbol{x},t)]=g(\boldsymbol{x},t),\quad(\boldsymbol{x},t)\in\partial\Omega,\end{gathered}$ (1)
@@
$\begin{split}\mathcal{L}_{\text{PINNs}}&=\lambda_{r}\sum_{i=1}^{N_{r}}\|\mathcal{D}[\hat{u}(\boldsymbol{x}_{i},t_{i})]-f(\boldsymbol{x}_{i},t_{i})\|^{2}+\lambda_{b}\sum_{i=1}^{N_{b}}\|\mathcal{B}[\hat{u}(\boldsymbol{x}_{i},t_{i})]-g(\boldsymbol{x}_{i},t_{i})\|^{2},\end{split}$ (2)
where $\{(\boldsymbol{x}_{i},t_{i})\}\subset\Omega$ are residual points, $\{(\boldsymbol{x}_{i},t_{i})\}\subset\partial\Omega$ are boundary/initial constraints or available data samples, and $\lambda_{r}$, $\lambda_{b}$ balance the loss terms. The goal is to train $\hat{u}$ to minimize $\mathcal{L}_{\text{PINNs}}$ using machine learning techniques.
@@ 3.1 Chebyshev1-Based Kolmogorov-Arnold Network Layer
Unlike traditional Kolmogorov-Arnold Networks (KAN) that employ spline coefficients, the First-kind Chebyshev KAN Layer leverages the properties of mesh-free Chebyshev polynomials to enhance both computational efficiency and approximation accuracy (chebykan2024, ; shukla2024comprehensive, ).
@@
To ensure the input values fall within the domain $[-1,1]$, Cheby1KAN applies the hyperbolic tangent function for normalization:
@@
[…]
@@
Compared to the original B-spline-based KANs, Chebyshev polynomials of the first kind in Equation (3) concentrate spectral energy in high frequencies with frequencies that increase linearly with the polynomial order $n$ (xu2024chebyshev, ; xiaoamortized, ), while maintaining global orthogonality over the interval $[-1,1]$:
[…]
@@ 3.2 Rank Diminution in Cheby1KAN Networks
While Cheby1KAN layers offer significant advantages, networks composed solely of stacked Cheby1KAN layers, as presented in Equation (7), exhibit pronounced rank diminution (feng2022rank, ). Consequently, these networks suffer a reduced capacity for feature representation, leading to severe information degradation and loss. We present a detailed derivation and proof of this phenomenon below (roth2024rank, ). The complete mathematical derivations are provided in Appendix B.
@@
Theorem 3.1 (Single Cheb1KAN Layer Rank Constraint).
@@
Theorem 3.2 (Nonlinear Normalization Effect).
@@
Theorem 3.3 (Exponential Decay in Infinite Depth).
@@
In summary, Cheby1KAN networks inherently experience rank diminution due to various factors. Collectively, the bounded rank per Cheby1KAN layer (Theorem 3.1), the attenuation from $\tanh(\cdot)$ (Theorem 3.2), and the multiplicative rank bound culminate in exponential rank decay (Theorem 3.3), thereby demonstrating the inherent rank diminution in Cheby1KAN networks.
Therefore, there is a significant need to improve the internal structure of models based on the Chebyshev1KAN layer, which will be discussed in detail in Section 3.3. Additionally, to address some computational limitations associated with the use of Cheby1KAN, we propose an external attention mechanism, which will be elaborated in Section 3.4. By incorporating both internal and external attention mechanisms, our AC-PKAN model fully leverages the advantages of Chebyshev Type-I polynomials while overcoming their initial drawbacks.
@@ 3.3 Internal Model Architecture
To resolve the Rank Diminution issue arising from direct stacking of Cheby1KAN layers in network architectures, we propose the AC-PKAN model, featuring an attention-enhanced framework (wang2021understanding, ; wang2024piratenets, ) designed to mitigate feature space collapse. The architecture synergistically combines linear transformations for input-output dimensional modulation, state-of-the-art activation functions, and residual-augmented Cheby1KAN layers. These components are collectively designed to preserve hierarchical feature diversity while capturing high-order nonlinear interactions and multiscale topological dependencies inherent in complex data structures. The algorithm’s details are provided in Algorithm 2 in Appendix A.
@@
To modulate the dimensionality of the data, the model employs linear transformations at both the input and output stages. The linear layer is designed to achieve a hybridization of KAN and MLP architectures. Its role as both an initial and final projection is inspired by the Spatio-Temporal Mixer linear layer in the PINNsformer model (zhao2023pinnsformer, ), which enhances spatiotemporal aggregation. The input features $\mathbf{x}$ are projected into a higher-dimensional space, and the final network representation $\alpha^{(L)}$ is mapped to the output space via:
@@
We adopt the state-of-the-art Wavelet activation function in the field of PINNs, as detailed in (zhao2023pinnsformer, ). Inspired by Fourier transforms, it introduces non-linearity and effectively captures periodic patterns:
@@
where $w_{1}$ and $w_{2}$ are learnable parameters initialized to one. This activation integrates Fourier feature embedding (wang2023expert, ) and sine activation (wong2022learning, ). When applied to encoders $U$ and $V$, the Wavelet activation preserves the gradient benefits introduced by the triangular activation function while modulating its phase and magnitude. This enhancement boosts representational capacity and facilitates adaptive Fourier embedding, thereby more effectively capturing periodic features and mitigating spectral bias.
@@
$\mathbf{U}=\text{Wavelet}(\mathbf{h}_{0}\mathbf{\Theta}_{U}+\mathbf{b}_{U}),\quad\mathbf{V}$ $=\text{Wavelet}(\mathbf{h}_{0}\mathbf{\Theta}_{V}+\mathbf{b}_{V}),$ (15)
@@
Our AC-PKAN’s inherent attention mechanism eliminates the need for an additional bias function $b(x)$ required in previous KAN models to maintain non-zero higher-order derivatives (wang2024kolmogorov, ). This reduces model complexity and parameter count while preserving the ability to seamlessly approximate PDEs of arbitrary finite order. By ensuring non-zero derivatives of any finite order and invoking the Kolmogorov–Arnold representation theorem, our model can approximate such PDEs.
Proposition 3.4.
@@
Proposition 3.5.
Let $\mathcal{N}$ be an AC-PKAN model with $L$ layers ($L\geq 2$) and infinite width. Then, the Jacobian matrix $J_{\mathcal{N}}(\boldsymbol{x})=\left[\frac{\partial\mathcal{N}_{i}}{\partial x_{j}}\right]_{m\times d}$ is full rank in the input space $\mathbb{R}^{d}$.
This property effectively addresses the internal rank diminution issue of Cheby1KAN networks discussed in Section 3.2, and also ensures stable gradient backpropagation, thereby preventing rank-deficiency-induced training failures in AC-PKAN.The complete mathematical derivations are provided in Appendix B.

### v2

@@ Abstract
Kolmogorov–Arnold Networks (KANs) have recently shown promise for solving partial differential equations (PDEs). Yet their original formulation is computationally and memory intensive, motivating the introduction of Chebyshev Type-I-based KANs (Cheby1KANs). Although Cheby1KANs have outperformed the vanilla KANs architecture, our rigorous theoretical analysis reveals that they still suffer from rank collapse, ultimately limiting their expressive capacity. To overcome these limitations, we enhance Cheby1KANs by integrating wavelet-activated MLPs with learnable parameters and an internal attention mechanism. We prove that this design preserves a full-rank Jacobian and is capable of approximating solutions to PDEs of arbitrary order. Furthermore, to alleviate the loss instability and imbalance introduced by the Chebyshev polynomial basis, we externally incorporate a Residual Gradient Attention (RGA) mechanism that dynamically re-weights individual loss terms according to their gradient norms and residual magnitudes. By jointly leveraging internal and external attention, we present AC-PKAN, a novel architecture that constitutes an enhancement to weakly supervised Physics-Informed Neural Networks (PINNs) and extends the expressive power of KANs. Experimental results from nine benchmark tasks across three domains show that AC-PKAN outperforms or matches state-of-the-art models such as PINNsFormer, establishing it as a highly effective tool for solving complex real-world engineering problems in zero-data or data-sparse regimes. The code is publicly available at https://github.com/fogradio/ACPKAN.
@@ 1 Introduction
Numerical solutions of partial differential equations (PDEs) are essential in science and engineering (Zienkiewicz and Taylor, 2005; Liu, 2009; Fornberg, 1998; Brebbia et al., 2012). Physics-informed neural networks (PINNs) (Lagaris et al., 1998; Raissi et al., 2019) have emerged as a promising approach in scientific machine learning (SciML), especially when data are unavailable or scarce. Traditional PINNs typically employ multilayer perceptrons (MLPs) (Cybenko, 1989) due to their ability to approximate nonlinear functions (Hornik et al., 1989) and their success in various PDE-solving applications (Yu and others, 2018; Han et al., 2018).
@@
However, PINNs encounter limitations, including difficulties with multi-scale phenomena (Kharazmi et al., 2021), the curse of dimensionality in high-dimensional spaces (Jagtap and Karniadakis, 2020), and challenges with nonlinear PDEs (Yuan et al., 2022). These issues arise from both the complexity of PDEs and limitations in PINN architectures and training methods. To address these challenges, existing methods focus on improving both the internal architecture of PINNs and their external learning strategies. Internal improvements include novel architectures like Quadratic Residual Networks (Qres) (Bu and Karpatne, 2021), First-Layer Sine (FLS) (Wong et al., 2022), and PINNsformer (Zhao et al., 2023). External strategies are discussed in detail in Section 2. Nevertheless, traditional PINNs based on MLPs still suffer from issues like lack of interpretability (Cranmer, 2023), overfitting, vanishing or exploding gradients, and scalability problems (Bachmann et al., 2024). As an alternative, Kolmogorov–Arnold Networks (KANs) (Liu et al., 2024b), inspired by the Kolmogorov–Arnold representation theorem (Kolmogorov, 1961; Braun and Griebel, 2009), have been proposed to offer greater accuracy and interpretability. KANs can be viewed as a combination of Kolmogorov networks and MLPs with learnable activation functions (Köppen, 2002; Sprecher and Draghici, 2002). Various KAN variants have emerged by replacing the B-spline functions (SS, 2024; Bozorgasl and Chen, 2024; Xu et al., 2024a). Although they still face challenges (Yu et al., 2024), KANs have shown promise in addressing issues like interpretability (Liu et al., 2024a) and catastrophic forgetting (Vaca-Rubio et al., 2024) in learning tasks (Samadi et al., 2024). Recent architectures like KINN (Wang et al., 2024b) and DeepOKAN (Abueidda et al., 2024) have applied KANs to PDE solving with promising results.
@@
Despite the potential of KANs, the original KAN suffers from high memory consumption and long training times due to the use of B-spline functions (Shukla et al., 2024). To address these limitations, we propose the Attention-Enhanced and Chebyshev Polynomial-Based Physics-Informed Kolmogorov–Arnold Networks (AC-PKAN). Our approach replaces B-spline functions with first-kind Chebyshev polynomials, forming the Cheby1KAN layer (SS, 2024), eliminating the need for grid storage and updates. Nevertheless, networks composed solely of stacked Cheby1KAN layers exhibit pronounced rank diminution (Feng et al., 2022). By integrating Cheby1KAN with linear layers and incorporating internal attention mechanisms derived from input features, AC-PKAN addresses these limitations while efficiently modeling complex nonlinear functions and selectively emphasizing distinct aspects of the input features at each layer. Additionally, we introduce an external attention mechanism that adaptively reweights loss terms according to both gradient norms and point-wise residuals, thereby counteracting the large polynomial expansions and gradient magnitudes inherent in Cheby1KAN, mitigating residual imbalance and gradient flow stiffness, and ultimately enhancing training stability and efficiency. To our knowledge, AC-PKAN is the first PINN framework to integrate internal and external attention mechanisms into KAN layers, effectively addressing many issues of original KANs and PINNs. Our key contributions can be summarized as follows:
@@ 2 Related Works
External Learning Strategies for PINNs. Most advances in physics-informed neural networks improve training from the outside while keeping an MLP-like backbone. Loss rebalancing methods such as PINN-LRA, PINN-NTK and residual-based adaptation (Wang et al., 2021; 2022b; Anagnostopoulos et al., 2024) adjust PDE, boundary and initial terms with gradient or NTK statistics to correct the mismatch among losses. Sampling based approaches follow the same goal. AAS selects collocation points using adversarial optimal transport (Tang et al., 2023), RoPINN applies regional Monte Carlo sampling (Pan et al., 2024), while RAR and PINNACLE resample high-residual areas and co-optimise all point types (Wu et al., 2023; Lau et al., 2024). Other works modify the objective or the training domain. gPINN adds gradient terms to enforce the PDE (Yu et al., 2022), vPINN uses a variational form (Kharazmi et al., 2019), LAAF and GAAF tune activations during training to accelerate convergence (Jagtap et al., 2020a; b), and FBPINN and hp-VPINN decompose the domain to make multi-scale problems tractable (Moseley et al., 2023; Kharazmi et al., 2021). These methods stabilise PINNs but they still rely on feature spaces that can degenerate when several operators or high order derivatives are imposed. Our approach keeps these mature external pipelines while adding an internal mechanism whose role is to preserve expressive bases during physics-informed optimisation.
KAN and Chebyshev-based Variants. Kolmogorov–Arnold Networks make the activation on edges learnable and can approximate operators with smaller models (Liu et al., 2024b). Follow up designs replace splines with faster or more structured bases, such as FastKAN with RBFs (Li, 2024), Cheby1KAN and Cheby2KAN with first and second kind polynomials for oscillatory targets (SS, 2024), rKAN and fKAN with rational or fractional Jacobi bases (Aghaei, 2024; 2025), and FourierKAN with Fourier modes (Mehrabian et al., 2024). Surveys place these models in a wider landscape of Kolmogorov-inspired approximators (Guilhoto and Perdikaris, 2024), and preliminary benchmarks still report Chebyshev-based variants as a strong speed–accuracy choice (SS, 2024). Yet most of these works optimise for approximation, interpretability or inference cost and do not discuss how to keep KANs stable when trained with collocation-based PDE losses. AC-PKAN is positioned at this intersection. It retains the approximation benefits of Chebyshev KANs, but augments them with internal feature re-injection, frequency-aware activation and rank-aware gating so that the model remains expressive under the same loss reweighting and sampling strategies used by advanced PINNs.
@@ 3 Motivation and Methodology
$\begin{gathered}\mathcal{D}[u(\bm{x},t)]=f(\bm{x},t),\quad(\bm{x},t)\in\Omega,\\ \mathcal{B}[u(\bm{x},t)]=g(\bm{x},t),\quad(\bm{x},t)\in\partial\Omega,\end{gathered}$ (1)
@@
$\begin{split}\mathcal{L}_{\text{PINNs}}&=\lambda_{r}\sum_{i=1}^{N_{r}}\|\mathcal{D}[\hat{u}(\bm{x}_{i},t_{i})]-f(\bm{x}_{i},t_{i})\|^{2}+\lambda_{b}\sum_{i=1}^{N_{b}}\|\mathcal{B}[\hat{u}(\bm{x}_{i},t_{i})]-g(\bm{x}_{i},t_{i})\|^{2},\end{split}$ (2)
where $\{(\bm{x}_{i},t_{i})\}\subset\Omega$ are residual points, $\{(\bm{x}_{i},t_{i})\}\subset\partial\Omega$ are boundary/initial constraints or available data samples, and $\lambda_{r}$, $\lambda_{b}$ balance the loss terms. The goal is to train $\hat{u}$ to minimize $\mathcal{L}_{\text{PINNs}}$ using machine learning techniques.
@@ 3.1 Chebyshev1-Based Kolmogorov-Arnold Network Layer
Unlike traditional Kolmogorov-Arnold Networks (KAN) that employ spline coefficients, the First-kind Chebyshev KAN Layer leverages the properties of mesh-free Chebyshev polynomials to enhance both computational efficiency and approximation accuracy (SS, 2024; Shukla et al., 2024).
@@
To ensure the input values fall within the domain [1,1][-1,1], Cheby1KAN applies the hyperbolic tangent function for normalization:
@@
${\rm Cheby1KAN}(\mathbf{x})=(\mathbf{\Phi}_{L-1}\circ\mathbf{\Phi}_{L-2}\circ\cdots\circ\mathbf{\Phi}_{1}\circ\mathbf{\Phi}_{0})\mathbf{x}.$ (8)
@@
Compared to the original B-spline-based KANs, Chebyshev polynomials of the first kind in Equation (3) concentrate spectral energy in high frequencies with frequencies that increase linearly with the polynomial order $n$ (Xu et al., 2024b; Xiao et al., 2024), while maintaining global orthogonality over the interval [1,1][-1,1]:
$\begin{split}\int_{-1}^{1}\frac{T_{m}(x)T_{n}(x)}{\sqrt{1-x^{2}}}\,dx=\begin{cases}0&m\neq n,\\ \pi&m=n=0,\\ \pi/2&m=n\neq 0.\end{cases}\end{split}$ (9)
This global support and slower decay of high-frequency components outperform locally supported B-splines, which lack global orthogonality and have rapidly diminishing high-frequency capture. Furthermore, Cheby1KAN layers require only a coefficient matrix of size $(\text{input\_dim},\text{output\_dim},\text{degree}+1)$, whereas B-spline-based KANs necessitate storing grids of size $(\text{in\_features},\text{grid\_size}+2\times\text{spline\_order}+1)$ and coefficient matrices of size $(\text{out\_features},\text{in\_features},\text{grid\_size}+\text{spline\_order})$, in addition to generating polynomial bases, solving local interpolation systems, and performing recursive updates to achieve high-order interpolation within their support intervals (Liu et al., 2024b). Hence, the Cheby1KAN layer significantly reduces both computational and memory overhead compared to the original B-spline-based KANs, while more effectively capturing high-frequency features. More details can be found at Appendix B
@@ 3.2 Rank Diminution in Cheby1KAN Networks
While Cheby1KAN layers offer significant advantages, networks composed solely of stacked Cheby1KAN layers, as presented in equation 8, exhibit pronounced rank diminution (Feng et al., 2022). Consequently, these networks suffer a reduced capacity for feature representation, leading to severe information degradation and loss. We present a detailed derivation and proof of this phenomenon below (Roth and Liebig, 2024). The complete mathematical derivations are provided in Appendix A.
@@
Theorem 1 (Single Cheb1KAN Layer Rank Constraint).
@@
Theorem 2 (Nonlinear Normalization Effect).
@@
Theorem 3 (Exponential Decay in Infinite Depth).
@@
In summary, Cheby1KAN networks inherently experience rank diminution due to various factors. Collectively, the bounded rank per Cheby1KAN layer (Theorem 1), the attenuation from $\tanh(\cdot)$ (Theorem 2), and the multiplicative rank bound culminate in exponential rank decay (Theorem 3), thereby demonstrating the inherent rank diminution in Cheby1KAN networks.
Therefore, there is a significant need to improve the internal structure of models based on the Cheby1KAN layer, which will be discussed in detail in Section 3.3. Additionally, to address some computational limitations associated with the use of Cheby1KAN, we propose an external attention mechanism, which will be elaborated in Section 3.4. By incorporating both internal and external attention mechanisms, our AC-PKAN model fully leverages the advantages of Chebyshev Type-I polynomials while overcoming their initial drawbacks.
@@ 3.3 Internal Model Architecture
To resolve the Rank Diminution issue arising from direct stacking of Cheby1KAN layers in network architectures, we propose the AC-PKAN model, featuring an attention-enhanced framework (Wang et al., 2021; 2024a) designed to mitigate feature space collapse. The architecture synergistically combines linear transformations for input-output dimensional modulation, state-of-the-art activation functions, and residual-augmented Cheby1KAN layers. These components are collectively designed to preserve hierarchical feature diversity while capturing high-order nonlinear interactions and multiscale topological dependencies inherent in complex data structures. The algorithm’s details are provided in Algorithm 1.
@@
To modulate the dimensionality of the data, the model employs linear transformations at both the input and output stages. The linear layer is designed to achieve a hybridization of KAN and MLP architectures. Its role as both an initial and final projection is inspired by the Spatio-Temporal Mixer linear layer in the PINNsformer model (Zhao et al., 2023), which enhances spatiotemporal aggregation. The input features $\mathbf{x}$ are projected into a higher-dimensional space, and the final network representation $\alpha^{(L)}$ is mapped to the output space via:
@@
We adopt the state-of-the-art Wavelet activation function in the field of PINNs, as detailed in (Zhao et al., 2023). Inspired by Fourier transforms, it introduces non-linearity and effectively captures periodic patterns:
@@
where $w_{1}$ and $w_{2}$ are learnable parameters initialized to one. This activation integrates Fourier feature embedding (Wang et al., 2023) and sine activation (Wong et al., 2022). When applied to encoders $U$ and $V$, the Wavelet activation preserves the gradient benefits introduced by the triangular activation function while modulating its phase and magnitude. This enhancement boosts representational capacity and facilitates adaptive Fourier embedding, thereby more effectively capturing periodic features and mitigating spectral bias.
@@
$\mathbf{U}=\text{Wavelet}(\mathbf{h}_{0}\mathbf{\Theta}_{U}+\mathbf{b}_{U}),\qquad\mathbf{V}$ $=\text{Wavelet}(\mathbf{h}_{0}\mathbf{\Theta}_{V}+\mathbf{b}_{V}),$ (16)
@@
Algorithm 1 Internal AC-PKAN Forward Pass
Data: Input data $\mathbf{x}$, Cheby1KAN layer parameters, Wavelet activation parameters
1:Input embedding:
2:Compute representations:
$\mathbf{U}\leftarrow\mathrm{Wavelet}(\mathbf{h}_{0}\mathbf{\Theta}_{U}+\mathbf{b}_{U}),\quad\mathbf{V}\leftarrow\mathrm{Wavelet}(\mathbf{h}_{0}\mathbf{\Theta}_{V}+\mathbf{b}_{V})$
3:Initialize attention: $\alpha^{(0)}\leftarrow\mathbf{U}$
4:for $l=1$ to $L$ do
5:  $\mathbf{H}^{(l)}\leftarrow\mathrm{LayerNorm}\!\big(\mathrm{Cheby1KANLayer}(\alpha^{(l-1)})\big)$
6:  $\alpha^{(l)}_{0}\leftarrow\mathbf{H}^{(l)}+\alpha^{(l-1)}$
7:  $\alpha^{(l)}\leftarrow(1-\alpha^{(l)}_{0})\odot\mathbf{U}+\alpha^{(l)}_{0}\odot(\mathbf{V}+1)$
8:end for
9:Output prediction:
@@
Our AC-PKAN’s inherent attention mechanism eliminates the need for an additional bias function $b(x)$ required in previous KAN models to maintain non-zero higher-order derivatives (Wang et al., 2024b). This reduces model complexity and parameter count while preserving the ability to seamlessly approximate PDEs of arbitrary finite order. By ensuring non-zero derivatives of any finite order and invoking the Kolmogorov–Arnold representation theorem, our model can approximate such PDEs.
Proposition 1.
@@
Proposition 2.
Let $\mathcal{N}$ be an AC-PKAN model with $L$ layers ($L\geq 2$) and infinite width. Then, the Jacobian matrix $J_{\mathcal{N}}(\bm{x})=\left[\frac{\partial\mathcal{N}_{i}}{\partial x_{j}}\right]_{m\times d}$ is full rank in the input space $\mathbb{R}^{d}$.
This property effectively addresses the internal rank diminution issue of Cheby1KAN networks discussed in Section 3.2, and also ensures stable gradient backpropagation, thereby preventing rank-deficiency-induced training failures in AC-PKAN.The complete mathematical derivations are provided in Appendix A.